22 MEN CAN COMPLETE A PIECE OF WORK IN 17 DAYS. THEY WORK FOR 2 DAYS. HOW MANY MORE MEN SHOULD NOW BE EMPLOYED SO AS TO COMPLETE THE WORK IN ANOTHER 10 DAYS?

Grade:8

1 Answers

Nitesh
6 years ago
Answer is 11 more men required.work done by 22 men in two days is 2/17.Remaining work is 1-2/17=15/17.This 15/17 work have to be done in 10 days.So from unitary method we get 22×15/10=33.So the extra men needed is 33-22=11.hence answer is 11